OpenStudy (anonymous):

ok I need to know if this a direct variation relationship? x= 3 -1 -4 y=-6 2 8

OpenStudy (jdoe0001):

\( \begin{array}{rrrllll} x&y\\ \hline\\ 3&-6&\implies 3\cdot {\color{red}{ -2}}\\ -1&2&\implies -1\cdot {\color{red}{ -2}}\\ 4&8&\implies -4\cdot {\color{red}{ -2}} \end{array}\)

OpenStudy (jdoe0001):

\(\begin{array}{llll} \textit{something }&\textit{varies directly to }&\textit{something else}\\ \quad \\ \textit{something }&=(\textit{some value })&\textit{something else}\\ \quad \\ y&=(n)&x \end{array}\)